The greatest number which always divides the product of the predecessor and successor of an odd natural number other than 1, is 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume an odd natural number be 15.
Then, predecessor of 5 = 5 -1 = 14
Successor of 5 = 5 + 1 =16
Then, the product of predecessor and successor = 14 × 16

= 224

24 is divided by 4 = 224/4 =56.
similarly, we can see and observe the pattern and we will get 4 as that number
Thus, the greatest number which always divides the product of the predecessor and successor of an odd natural number other than 1, is 4.

Thanks